OpenStudy (anonymous):

4x^-2 y^8 z^0 / 3x y z^4 simplify

OpenStudy (anonymous):

\[\dfrac{4x^{-2}y^8z^0}{3xyz^4}=\dfrac{4}{3}\cdot\dfrac{x^{-2}}{x}\cdot \dfrac{y^8}{y}\cdot\dfrac{z^0}{4}=\dfrac{4}{3}x^{-2-1}y^{8-1}z^{0-4}=\dfrac{4}{3}x^{-3}y^7z^{-4}\]This is becuase:\[\dfrac{x^a}{x^b}=x^{a-b}\]